ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 222, 2017

Aistleitner, C., G. Larcher and M. Lewko Additive energy and the Hausdorff dimension of the exceptional set in metric pair correlation problems

463

Alexander, M., M. Henk and A. Zvavitch A discrete version of Koldobsky’s slicing inequality

261

Alon, N., A. Pokrovskiy and B. Sudakov Random subgraphs of properly edge-coloured complete graphs and long rainbow cycles

317

Bandle, C., M. Marcus and V. Moroz Boundary singularities of solutions of semilinear elliptic equations in the half-space with a Hardy potential

487

Baztokos, S., S. Dann, A. Giannopoulos and A. Koldbosky On the average volume of sections of convex bodies

921

Bédaride, N. and T. Fernique Weak local rules for planar octagonal tilings

63

Bergeron, N. and L. Clozel Sur la cohomologie des variétés hyperboliques de dimension 7 trialitaires

333

Billig, Y. and V. Futorny Classification of simple cuspidal modules for solenoidal Lie algebras

109

Binder, J. Fields of rationality of cusp forms  973

Bleak, C., C. Donoven and J. Jonušas Some isomorphism results for Thompson-like groups V_{n}(G)

1

Boralevi, A., J. Draisma, E. Horobeţ and E. Robeva Orthogonal and unitary tensor decomposition from an algebraic perspective

223

Boshernitzan, M., N. Frantzikinakis and M. Wierdl Underrecurrence in the Khintchine recurrence theorem

815

Bukh, B. Ranks of matrices with few distinct entries  165

Bureaux, J. and N. Enriquez Asymptotics of convex lattice polygonal lines with a constrained number of vertices

515

Castillo, J. M. F., W. Cuellar, V. Ferenczi and Y. Moreno Complex structures on twisted Hilbert spaces

787

Chang, M.-C. Arithmetic progressions in multiplicative groups of finite fields

631

Clozel, L. See N. Bergeron Cuellar, W. See J. M. F. Castillo Dann, S. See S. Baztokos de Pagter, B., P. G. Dodds and F. A. Sukochev On weak^{∗}

convergent sequences in duals of symmetric spaces of τ-measurable operators

125

Dodds, P. G. See B. de Pagter Donoven, C. See C. Bleak Draisma, J. See A. Boralevi Dykema, K., F. Sukochev and D. Zanin An upper triangular de-

composition theorem for some unbounded operators affiliated to II_{1}- factors

645

Eleftherakis, G. K., E. T. Kakariadis and E. G. Katsoulis Morita equivalence of C*-correspondences passes to the related operator algebras

949

Enriquez, N. See J. Bureaux Ferenczi, V. See J. M. F. Castillo Fernique, T. See N. Bédaride

Frankl, P. Proof of the Erdős matching conjecture in a new range  421

Frantzikinakis, N. See M. Boshernitzan Futorny, V. See Y. Billig Giannopoulos, A. See S. Baztokos Glebov, R., Z. Luria and B. Sudakov The number of Hamiltonian

decompositions of regular graphs

91

Goaoc, X., I. Mabillard, P. Paták, Z. Patáková, M. Tancer and U. Wagner On generalized Heawood inequalities for manifolds: a van Kampen–Flores-type nonembeddability result

841

Goertsches, O., H. Nozawa and D. Töben Localization of Chern– Simons type invariants of Riemannian foliations

867

Henk, M. See M. Alexander Horobeţ, E. See A. Boralevi Jahnel, J. and D. Schindler On integral points on degree four del

Pezzo surfaces

21

Jonušas, J. See C. Bleak Kakariadis, E. T. See G. K. Eleftherakis Kaplan, E. The characterization of theta-distinguished representa-

tions of GL(n)

551

Katsoulis, E. G. See G. K. Eleftherakis Khukhro, E. I. and P. Shumyatsky Engel-type subgroups and

length parameters of finite groups

599

Koldbosky, A. See S. Baztokos Larcher, G. See C. Aistleitner Lee, S. H., W. Y. Lee and J. Yoon An answer to a question of

A. Lubin: The lifting problem for commuting subnormals

201

Lee, W. Y. See S. H. Lee Lewko, M. See C. Aistleitner Luria, Z. See R. Glebov Mabillard, I. See X. Goaoc Marcus, M. See C. Bandle Moreno, Y. See J. M. F. Castillo Moroz, V. See C. Bandle

Murai, S. and I. Novik Face numbers and the fundamental group  297

Novik, I. See S. Murai Nozawa, H. See O. Goertsches Paták, P. See X. Goaoc Patáková, Z. See X. Goaoc Pokrovskiy, A. See N. Alon Robeva, E. See A. Boralevi Schindler, D. See J. Jahnel Shumyatsky, P. See E. I. Khukrho Soudry, D. The unramified computation of Rankin–Selberg integrals

expressed in terms of Bessel models for split orthogonal groups: Part I

711

Sudakov, B. See N. Alon Sudakov, B. See R. Glebov Sukochev, F. A. See B. de Pagter Sukochev, F. See K. Dykema Tancer, M. See X. Goaoc Töben, D. See O. Goertsches Wagner, U. See X. Goaoc Wang, K. and G. Zhang Curvature inequalities for operators of the

Cowen–Douglas class

279

Westrick, L. B. Seas of squares with sizes from a Π_{1}^{0} set  431

Wierdl, M. See M. Boshernitzan Yoon, J. See S. H. Lee Zanin, D. See K. Dykema Zhang, G. See K. Wang

Zheng, H. The flag upper bound theorem for 3- and 5-manifolds  401

Zvavitch, A. See M. Alexander